A resistor of resistance 100 ohm is connected to an AC source E = (12V)sin (250πs^-1)t. Find the energy dissipated as heat during t=0 to t=1ms
1
Expert's answer
2020-07-27T13:52:13-0400

Heat energy dissipated "H=\int_0^{t}\space\frac{E^2}{R}dt"



where "t=10^{-3}", "E=E_{rms}=E_0sin\omega t" ,

Therefore


"\int_0^{t}\space \frac{(12sin\omega t)^{2}}{100}dt"


"\frac{144}{100}\int_0^{t}\space\frac{(1-cos2\times250\pi t)}{2}dt"


"\frac{144}{200}[1-\frac{sin500\pi t}{500\pi}]_0^t"


At "t=10^{-3}"

Energy dissipated ="\frac{144}{200}[10^{-3}-\frac{1}{500\pi}]"


Heat Energy= "2.61\times10^{-4}J"